# A Survey on 3D Gaussian Splatting

### v2

@@ 1 Introduction
The journey of 3D scene reconstruction began long before the surge of deep learning, with early endeavors focusing on light fields and basic scene reconstruction methods [1, 2, 3]. These early attempts, however, were limited by their reliance on dense sampling and structured capture, leading to significant challenges in handling complex scenes and lighting conditions. The emergence of structure-from-motion [4] and subsequent advancements in multi-view stereo [5] algorithms provided a more robust framework for 3D scene reconstruction. Despite these advancements, such methods struggled with novel-view synthesis and lacked compatibility with deep scene understanding models. NeRF represents a quantum leap in this progression. By leveraging deep neural networks, NeRF enabled the direct mapping of spatial coordinates to color and density. The success of NeRF hinged on its ability to create continuous, volumetric scene functions, producing results with unprecedented detail and realism. However, as with any burgeoning technology, this implementation came at a cost: i) Computational Intensity. NeRF based methods are computationally intensive [6, 7, 8, 9, 10, 11], often requiring extensive training times and substantial resources for rendering, especially for high-resolution outputs. ii) Editability. Manipulating scenes represented implicitly can be challenging, as direct modifications to the neural network’s weights are not intuitively related to changes in geometric or appearance properties of the scene.
@@
It is in this context that 3D Gaussian splatting (GS) [12] emerges, not merely as an incremental improvement but as a paradigm-shifting approach that redefines the boundaries of scene representation and rendering. While NeRF excelled in creating photorealistic images, the need for faster, more efficient rendering methods was becoming increasingly apparent, especially for applications that are highly sensitive to latency. 3D GS addressed this need by introducing an advanced, explicit scene representation that models a scene using millions of learnable 3D Gaussians in space. Unlike the implicit, coordinate-based models [13, 14, 15], 3D GS employs an explicit representation and highly parallelized workflows, facilitating more efficient computation and rendering. The innovation of 3D GS lies in its unique blend of the benefits of differentiable pipelines and point-based rendering techniques [16, 17, 18, 19, 20, 21]. By representing scenes with learnable 3D Gaussians, it preserves the strong fitting capability of continuous volumetric radiance fields, essential for high-quality image synthesis, while simultaneously avoiding the computational overhead associated with NeRF based methods (e.g., computationally expensive ray-marching, and unnecessary calculations in empty space).
The introduction of 3D GS is not just a technical advancement; it represents a fundamental shift in how we approach scene representation and rendering in computer vision and graphics. By enabling real-time rendering capabilities without compromising on visual quality, 3D GS opens up a plethora of possibilities for applications ranging from virtual reality and augmented reality to real-time cinematic rendering and beyond [22, 23, 24, 25]. This technology holds the promise of not only enhancing existing applications but also enabling new ones that were previously unfeasible due to computational constraints. Furthermore, 3D GS’s explicit scene representation offers unprecedented flexibility to control the objects and scene dynamics, a crucial factor in complex scenarios involving intricate geometries and varying lighting conditions [26, 27]. This level of editability, combined with the efficiency of the training and rendering process, positions 3D GS as a transformative force in shaping future developments in relevant fields.
@@
In an effort to assist readers in keeping pace with the swift evolution of 3D GS, we provide the first survey on 3D GS, which presents a systematic and timely collection of the most significant recent literature on the topic. Given that 3D GS is a very recent innovation (Fig. 1), our focus is on the diverse developments and contributions that have emerged since its introduction. The selected follow-up works of 3D GS are primarily sourced from top conferences. Our primary goal is to provide a thorough and up-to-date analysis of the theoretical foundations, remarkable developments, and burgeoning applications of 3D GS, highlighting its potential to revolutionize the field. Acknowledging the nascent yet rapidly evolving nature of 3D GS, this survey is inevitably a biased view, but we have strived to pinpoint and discuss the fundamental characteristics of influential works so as to encapsulate primary research trends.
A summary of the structure of this article can be found in Fig. 2, which is presented as follows: Sec. 2 provides a brief background on problem formulation, terminology, and related research domains. Sec. 3 introduces the essential insights of 3D GS, encompassing the novel-view synthesis with 3D Gaussians and the optimization nuances of 3D GS. Sec. 4 presents several fruitful directions that aim to improve the capabilities of the original 3D GS. Sec. 5 unveils the diverse application areas and tasks where 3D GS has made significant impacts, showcasing its versatility. Sec. 6 conducts performance comparison and analysis. Finally, Sec. 7 and 8 highlight the open questions for further research and conclude the survey. We aim to provide a balanced perspective, reflecting the current scope and future potential of 3D GS, and thereby serve as a valuable resource for researchers and practitioners eager to understand and contribute to this rapidly evolving domain.
@@ 2 Background
In this section, we first provide a brief formulation of radiance fields (Sec. 2.1), a key concept in scene rendering. It outlines two main types of radiance field representations: implicit, like NeRF [15], which uses neural networks for a straightforward yet computationally demanding rendering; and explicit, like grid [28], which employs discrete structures for quicker access but at the cost of higher memory use. Sec. 2.2 further establishes linkages with relevant fields such as scene reconstruction and rendering. For a comprehensive overview of radiance fields, scene reconstruction and representation, and rendering, please see the excellent surveys [29, 30, 31, 32, 33, 34].
@@ 2.1.1 Radiance Field
A radiance field is a representation of light distribution in a three-dimensional space, which captures how light interacts with surfaces and materials in the environment [30]. Mathematically, a radiance field can be described as a function $L\!:\!\mathbb{R}^{5}\!\mapsto\!\mathbb{R}^{+}$, where $L(x,y,z,\theta,\phi)$ maps a point in space $(x,y,z)$, and a direction specified by spherical coordinates $(\theta,\phi)$, to a non-negative radiance value. Radiance fields can be encapsulated through implicit or explicit representations, each with specific advantages for scene representation and rendering.
@@ 2.1.2 Implicit Radiance Field
This format allows for a differentiable and compact representation of complex scenes, albeit often at the cost of high computational load during rendering due to volumetric ray marching [12].
@@ 2.1.3 Explicit Radiance Field
In contrast, an explicit radiance field directly represents the distribution of light in a discrete spatial structure, such as a voxel grid or a set of points [28, 37]. Each element in this structure stores the radiance information for its respective location in space. This approach allows for more direct and often faster access to radiance data but at the cost of higher memory usage and potentially lower resolution. A generic form for an explicit radiance field representation can be written as:
@@ 2.2 Context and Terminology
Several techniques and research disciplines possess a close relationship with 3D GS, which will be described briefly in the following sections.
@@ 2.2.1 Scene Reconstruction and Rendering
[…]
Figure 3: An illustration of the forward process of 3D GS (Sec. 3.1). (a) The splatting step projects 3D Gaussians into image space. (b) 3D GS divides the image into multiple non-overlapping patches, i.e., tiles. (c) 3D GS replicates the Gaussians which cover several tiles, assigning each copy an identifier, i.e., a tile ID. (d) By rendering the sorted Gaussians, we can obtain all pixels within the tile. Note that the computational workflows for pixels and tiles are independent and can be done in parallel. Best viewed in color.
@@ 2.2.2 Neural Rendering and Radiance Fields
Neural rendering integrates deep learning with traditional graphics techniques to create photorealistic images [42, 43]. Early attempts utilized convolutional neural networks for estimating blending weights [40] or texture-space solutions [44, 45]. As mentioned in Sec. 2.1.1, the radiance field represents a function that describes the amount of light traveling in every direction through every point in space. NeRFs [15, 11, 10] use neural networks, typically MLPs, to model the radiance fields, enabling detailed and realistic scene rendering.
@@ 2.2.3 Volumetric Representations and Ray-Marching
Volumetric representations model objects and scenes not just as surfaces but as volumes filled with materials or empty space [46]. This approach allows for a more accurate rendering of phenomena like fog, smoke, or translucent materials. Ray-marching is a technique used with volumetric representations to render images by incrementally tracing the path of light through a volume [13, 14]. NeRF [15] shares the same spirit of volumetric ray-marching and introduces importance sampling and positional encoding to improve the quality of synthesized images. While providing high-quality results, volumetric ray-marching is computationally expensive, motivating the search for more efficient methods like 3D GS.
@@ 2.2.4 Point-based Rendering
Point-based rendering is a technique for visualizing 3D scenes using points rather than traditional polygons. This method is particularly effective for rendering complex, unstructured, or sparse geometric data. Points can be augmented with additional properties like learnable neural descriptors [47, 48], and rendered efficiently [49, 50], but this approach suffers from issues like holes in the rendering or aliasing effects. 3D GS [12] extends this concept by using anisotropic Gaussians for a more continuous and cohesive representation of the scene. More details will be further discussed in Sec. 3.
@@
3D GS offers a breakthrough in real-time, high-resolution image rendering, without relying on neural components. This section aims to provide essential insights of 3D GS. We first elaborate on how 3D GS synthesizes an image given well-constructed 3D Gaussians in Sec. 3.1, i.e., the forward process of 3D GS. Then, we introduce how to obtain well-constructed 3D Gaussians for a given scene in Sec. 3.2, i.e., the optimization process of 3D GS.
3.1 Novel-view Synthesis with Learned 3D Gaussians
Consider a scene represented by (millions of) optimized 3D Gaussians. The objective is to generate an image from a specified camera pose. Recall that NeRFs approach this task through computationally demanding volumetric ray-marching, sampling 3D space points per pixel. Such a paradigm struggles with high-resolution image synthesis, failing to achieve real-time rendering, especially for platforms with limited computing resources [12]. By contrast, 3D GS begins by projecting these 3D Gaussians onto a pixel-based image plane, a process termed “splatting” (see Fig. 3a). Afterwards, 3D GS sorts these Gaussians and computes the value for each pixel. As shown in Fig. 4, the rendering of NeRFs and 3D GS can be viewed as an inverse process of each other. In what follows, we begin with the definition of a 3D Gaussian, which is the minimal element of the scene representation in 3D GS. Next, we describe how these 3D Gaussians can be used for differentiable rendering. Finally, we introduce the acceleration technique used in 3D GS, which is the key to fast rendering.
@@ 3.1 Novel-view Synthesis with Learned 3D Gaussians
$\bullet$ Splatting. In this step, 3D Gaussians (ellipsoids) are projected into the 2D image space (ellipses) for rendering. Given the viewing transformation $\bm{W}$ and 3D covariance matrix $\bm{\Sigma}$, the projected 2D covariance matrix $\bm{\Sigma}^{\prime}$ is computed using:
@@
$C=\sum_{i\in\mathcal{N}}c_{i}\alpha^{\prime}_{i}\prod_{j=1}^{i-1}\left(1-\alpha^{\prime}_{j}\right),$ (5)
@@
where $c_{i}$ is the learned color. The final opacity $\alpha^{\prime}_{i}$ is the multiplication result of the learned opacity $\alpha_{i}$ and the Gaussian, defined as follows:
$\alpha^{\prime}_{i}=\alpha_{i}\times\text{exp}\big{(}-\frac{1}{2}(\bm{x}^{\prime}-\bm{\mu}^{\prime}_{i})^{\top}\bm{\Sigma}^{\prime-1}_{i}(\bm{x}^{\prime}-\bm{\mu}^{\prime}_{i})\big{)},$ (6)
@@
where $\bm{x}^{\prime}$ and $\bm{\mu}^{\prime}_{i}$ are coordinates in the projected space. It is a reasonable concern that the rendering process described could be slower compared to NeRFs, given that generating the required sorted list is hard to parallelize. Indeed, this concern is justified; rendering speeds can be significantly impacted when utilizing such a simplistic, pixel-by-pixel approach. To achieve real-time rendering, 3D GS makes several concessions to accommodate parallel computation.
@@
$\bullet$ Tiles (Patches). To avoid the cost computation of deriving Gaussians for each pixel, 3D GS shifts the precision from pixel-level to patch-level detail. Concretely, 3D GS initially divides the image into multiple non-overlapping patches, called “tiles” in the original paper [12]. Fig. 3b provides an illustration of tiles. Each tile comprises 16$\times$16 pixels as suggested in [12]. 3D GS further determines which tiles intersect with these projected Gaussians. Given that a projected Gaussian may cover several tiles, a logical method involves replicating the Gaussian, assigning each copy an identifier (i.e., a tile ID) for the relevant tile.
@@
$\bullet$ Parallel Rendering. After replication, 3D GS combines the respective tile ID with the depth value obtained from the view transformation for each Gaussian. This results in an unsorted list of bytes where the upper bits represent the tile ID and the lower bits signify depth. By doing so, the sorted list can be directly utilized for rendering (i.e., alpha compositing). Fig. 3c and Fig. 3d provide the visual demonstration of such concepts. It’s worth highlighting that rendering each tile and pixel occurs independently, making this process highly suitable for parallel computations. An additional benefit is that each tile’s pixels can access a common shared memory and maintain a uniform read sequence, enabling parallel execution of alpha compositing with increased efficiency. In the official implementation of the original paper [12], the framework regards the processing of tiles and pixels as analogous to the blocks and threads, respectively, in CUDA programming architecture.
@@
In a nutshell, 3D GS introduces several approximations during the forward processing phase to enhance computational efficiency while maintaining a high standard of image synthesis quality.

### v3

@@ 1 Introduction
The journey of 3D scene reconstruction began long before the surge of deep learning, with early endeavors focusing on light fields and basic scene reconstruction methods [1, 2, 3]. These early attempts, however, were limited by their reliance on dense sampling and structured capture, leading to significant challenges in handling complex scenes and lighting conditions. The emergence of structure-from-motion [4] and subsequent advancements in multi-view stereo [5] algorithms provided a more robust framework for 3D scene reconstruction. Despite these advancements, such methods struggled with novel-view synthesis and lacked compatibility with deep scene understanding models. NeRF represents a quantum leap in this progression. By leveraging deep neural networks, NeRF enabled the direct mapping of spatial coordinates to color and density. The success of NeRF hinged on its ability to create continuous, volumetric scene functions, producing results with unprecedented detail and realism. However, as with any burgeoning technology, this implementation came at a cost: i) Computational Intensity. NeRF based methods are computationally intensive [6, 7, 8, 9], often requiring extensive training times and substantial resources for rendering, especially for high-resolution outputs. ii) Editability. Manipulating scenes represented implicitly is challenging, since direct modifications to the neural network’s weights are not intuitively related to changes in geometric or appearance properties of the scene.
@@
It is in this context that 3D Gaussian splatting (GS) [10] emerges, not merely as an incremental improvement but as a paradigm-shifting approach that redefines the boundaries of scene representation and rendering. While NeRF excelled in creating photorealistic images, the need for faster, more efficient rendering methods was becoming increasingly apparent, especially for applications (e.g., virtual reality and autonomous driving) that are highly sensitive to latency. 3D GS addressed this need by introducing an advanced, explicit scene representation that models a scene using millions of learnable 3D Gaussians in space. Unlike the implicit, coordinate-based models [11, 12], 3D GS employs an explicit representation and highly parallelized workflows, facilitating more efficient computation and rendering. The innovation of 3D GS lies in its unique blend of the benefits of differentiable pipelines and point-based rendering techniques [13, 14, 15, 16, 17]. By representing scenes with learnable 3D Gaussians, it preserves the strong fitting capability of continuous volumetric radiance fields, essential for high-quality image synthesis, while simultaneously avoiding the computational overhead associated with NeRF based methods (e.g., computationally expensive ray-marching, and unnecessary calculations in empty space).
The introduction of 3D GS is not just a technical advancement; it represents a fundamental shift in how we approach scene representation and rendering in computer vision and graphics. By enabling real-time rendering capabilities without compromising on visual quality, 3D GS opens up a plethora of possibilities for applications ranging from virtual reality and augmented reality to real-time cinematic rendering and beyond [18, 19, 20, 21]. This technology holds the promise of not only enhancing existing applications but also enabling new ones that were previously unfeasible due to computational constraints. Furthermore, 3D GS’s explicit scene representation offers unprecedented flexibility to control the objects and scene dynamics, a crucial factor in complex scenarios involving intricate geometries and varying lighting conditions [22, 23, 24]. This level of editability, combined with the efficiency of the training and rendering process, positions 3D GS as a transformative force in shaping future developments in relevant fields.
@@
In an effort to assist readers in keeping pace with the swift evolution of 3D GS, we provide the first survey on 3D GS, which presents a systematic and timely collection of the most significant literature on the topic. Given that 3D GS is a very recent innovation (Fig. 1), this survey focuses in particular on its principles, and the diverse developments and contributions that have emerged since its introduction. The selected follow-up works are primarily sourced from top-tier conferences, to provide a thorough and up-to-date analysis of the theoretical foundations, remarkable developments, and burgeoning applications of 3D GS. Acknowledging the nascent yet rapidly evolving nature of 3D GS, this survey is inevitably a biased view, but we strive to offer a balanced perspective that reflects both the current state and the future potential of this field. Our aim is to encapsulate the primary research trends and serve as a valuable resource for both researchers and practitioners eager to understand and contribute to this rapidly evolving domain.
A summary of the structure of this article can be found in Fig. 2, which is presented as follows: Sec. 2 provides a brief background on problem formulation, terminology, and related research domains. Sec. 3 introduces the essential insights of 3D GS, encompassing the rendering process with learned 3D Gaussians and the optimization details (i.e., how to learn 3D Gaussians) of 3D GS. Sec. 4 presents several fruitful directions that aim to improve the capabilities of the original 3D GS. Sec. 5 unveils the diverse application areas and tasks where 3D GS has made significant impacts, showcasing its versatility. Sec. 6 conducts performance comparison and analysis. Finally, Sec. 7 and 8 highlight the open questions for further research and conclude the survey.
@@ 2 Background
In this section, we first provide a brief formulation of radiance fields (Sec. 2.1), a key concept in scene rendering. It outlines two main types of radiance field representations: implicit, like NeRF [12], which uses neural networks for a straightforward yet computationally demanding rendering; and explicit, like grid [25], which employs discrete structures for quicker access but at the cost of higher memory use. Sec. 2.2 further establishes linkages with relevant fields such as scene reconstruction and rendering. For a comprehensive overview of radiance fields, scene reconstruction and representation, and rendering methods, please see the excellent surveys [26, 27, 28, 29, 30, 31] for more insights.
@@ 2.1.1 Radiance Field
A radiance field is a representation of light distribution in a 3D space, which captures how light interacts with surfaces and materials in the environment [27]. Mathematically, a radiance field can be described as a function $L\!:\!\mathbb{R}^{5}\!\mapsto\!\mathbb{R}^{+}$, where $L(x,y,z,\theta,\phi)$ maps a point in space $(x,y,z)$, and a direction specified by spherical coordinates $(\theta,\phi)$, to a non-negative radiance value. Radiance fields can be encapsulated through implicit or explicit representations, each with specific advantages for scene representation and rendering.
@@ 2.1.2 Implicit Radiance Field
This format allows for a differentiable and compact representation of complex scenes, albeit often at the cost of high computational load due to volumetric ray marching [10].
@@ 2.1.3 Explicit Radiance Field
In contrast, an explicit radiance field directly represents the distribution of light in a discrete spatial structure, such as a voxel grid or a set of points [25, 34]. Each element in this structure stores the radiance information for its respective location in space. This allows for more direct and often faster access to radiance data but at the cost of higher memory usage and potentially lower resolution. A generic form for an explicit radiance field representation can be written as:
@@ 2.2 Context and Terminology
Several techniques and research disciplines possess a close relationship with 3D GS, which will be described briefly in the following sections for clarity.
@@ 2.2.2 Neural Rendering and Radiance Fields
Neural rendering integrates deep learning with traditional graphics techniques to create photorealistic images. Early attempts utilized convolutional networks for estimating blending weights [37] or texture-space solutions [39]. As mentioned in Sec. 2.1.1, the radiance field represents a function that describes the amount of light traveling in every direction through every point in space. NeRFs [12, 9, 8] use neural networks, typically MLPs, to model the radiance fields, enabling detailed and realistic scene rendering.
@@ 2.2.3 Volumetric Representations and Ray-Marching
Volumetric representations model objects and scenes not just as surfaces but as volumes filled with materials or empty space. This allows for a more accurate rendering of phenomena like fog, smoke, or translucent materials. Ray-marching is a technique used with volumetric representations to render images by tracing the path of light through a volume [11]. NeRF [12] shares the same spirit of volumetric ray-marching and introduces importance sampling and positional encoding to improve the quality of synthesized images. While providing high-quality results, volumetric ray-marching is computationally expensive, motivating the search for more efficient methods like 3D GS.
@@ 2.2.4 Point-based Rendering
Point-based rendering is a technique for visualizing 3D scenes using points rather than traditional polygons. This method is particularly effective for rendering complex, unstructured, or sparse geometric data. Points can be augmented with additional properties like learnable neural descriptors [40, 41], and rendered efficiently [42, 43], but this approach suffers from issues like holes in the rendering or aliasing effects. 3D GS [10] extends this concept by using anisotropic Gaussians for a more continuous and cohesive representation of the scene. More implementation details will be further discussed in Sec. 3.
@@
3D GS offers a breakthrough in real-time, high-resolution image rendering, without relying on deep neural networks. This section aims to provide essential insights of 3D GS. We first elaborate on how 3D GS synthesizes an image given well-constructed 3D Gaussians in Sec. 3.1, i.e., the forward process of 3D GS. Then, we introduce how to obtain well-constructed 3D Gaussians for a given scene in Sec. 3.2, i.e., the optimization process of 3D GS.
3.1 Rendering with Learned 3D Gaussians
Consider a scene represented by (millions of) optimized 3D Gaussians. The objective is to generate an image from a specified camera pose. Recall that NeRFs approach this task through computationally demanding volumetric ray-marching, sampling 3D space points per pixel. Such a paradigm struggles with high-resolution image synthesis, failing to achieve real-time rendering, especially for platforms with limited computing resources [10]. By contrast, 3D GS begins by projecting these 3D Gaussians onto a pixel-based image plane, a process termed “splatting” (see Fig. 3b). Afterwards, 3D GS sorts these Gaussians and computes the value for each pixel. As shown in Fig. 3, the rendering of NeRFs and 3D GS can be viewed as an inverse process of each other. In what follows, we begin with the definition of a 3D Gaussian, which is the minimal element of the scene representation in 3D GS. Next, we describe how these 3D Gaussians can be used for differentiable rendering. Finally, we introduce the acceleration technique used in 3D GS, which is the key to fast rendering.
@@ 3.1 Rendering with Learned 3D Gaussians
$\bullet$ Splatting. In this step, 3D Gaussians (ellipsoids) in 3D space are projected into the 2D image space (ellipses) for rendering. Given the viewing transformation $\bm{W}$ and 3D covariance matrix $\bm{\Sigma}$, the projected 2D covariance matrix $\bm{\Sigma}^{\prime}$ is computed using:
@@
[…]
Figure 4: An illustration of the forward process of 3D GS (see Sec. 3.1). (a) The splatting step projects 3D Gaussians into image space. (b) 3D GS divides the image into multiple non-overlapping patches, i.e., tiles. (c) 3D GS replicates the Gaussians which cover several tiles, assigning each copy an identifier, i.e., a tile ID. (d) By rendering the sorted Gaussians, we can obtain all pixels within the tile. Note that the computational workflows for pixels and tiles are independent and can be done in parallel. Best viewed in color.
@@
$C=\sum_{n=1}^{|\mathcal{N}|}c_{n}\alpha^{\prime}_{n}\prod_{j=1}^{n-1}\left(1-\alpha^{\prime}_{j}\right),$ (5)
@@
where $c_{n}$ is the learned color. The final opacity $\alpha^{\prime}_{n}$ is the multiplication result of the learned opacity $\alpha_{n}$ and the Gaussian, defined as follows:
$\alpha^{\prime}_{n}=\alpha_{n}\times\text{exp}\big{(}-\frac{1}{2}(\bm{x}^{\prime}-\bm{\mu}^{\prime}_{n})^{\top}\bm{\Sigma}^{\prime-1}_{n}(\bm{x}^{\prime}-\bm{\mu}^{\prime}_{n})\big{)},$ (6)
@@
where $\bm{x}^{\prime}$ and $\bm{\mu}^{\prime}_{n}$ are coordinates in the projected space. It is a reasonable concern that the rendering process described could be slower compared to NeRFs, given that generating the required sorted list is hard to parallelize. Indeed, this concern is justified; rendering speeds can be significantly impacted when utilizing such a simplistic, pixel-by-pixel approach. To achieve real-time rendering, 3D GS makes several concessions to accommodate parallel computation.
5
Figure 5: An illustration of the tile based parallel (at the pixel-level) rendering. All the pixels within a tile (Tile1 here) access the same ordered Gaussian list stored in a shared memory for rendering. As the system processes each Gaussian sequentially, every pixel in the tile evaluates the Gaussian’s contribution according to the distance (i.e., the $\exp$ term in Eq. 6). Therefore, the rendering for a tile can be completed by iterating through the list of Gaussians just once. The computation for the red Gaussian follows a similar way and is omitted here for simplicity.
@@
$\bullet$ Tiles (Patches). To avoid the cost computation of deriving Gaussians for each pixel, 3D GS shifts the precision from pixel-level to patch-level detail. Concretely, 3D GS initially divides the image into multiple non-overlapping patches (called “tiles” in the original paper [10]). Fig. 4b provides an illustration of tiles. Each tile comprises 16$\times$16 pixels as suggested in [10]. 3D GS further determines which tiles intersect with these projected Gaussians. Given that a projected Gaussian may cover several tiles, a logical method involves replicating the Gaussian, assigning each copy an identifier (i.e., a tile ID) for the relevant tile.
@@
$\bullet$ Parallel Rendering. After replication, 3D GS combines the respective tile ID with the depth value obtained from the view transformation for each Gaussian. This results in an unsorted list of bytes where the upper bits represent the tile ID and the lower bits signify depth. By doing so, the sorted list can be directly utilized for rendering (i.e., alpha compositing). Fig. 4c and Fig. 4d provide the visual demonstration of such concepts. It’s worth highlighting that rendering each tile and pixel occurs independently, making this process highly suitable for parallel computations. An additional benefit is that each tile’s pixels can access a common shared memory and maintain an uniform read sequence (Fig. 5), enabling parallel execution of alpha compositing with increased efficiency. In the official implementation of the original paper [10], the framework regards the processing of tiles and pixels as analogous to the blocks and threads, respectively, in CUDA programming architecture.
@@
In a nutshell, 3D GS introduces several approximations during rendering to enhance computational efficiency while maintaining a high standard of image synthesis quality.